How many points with integer coordinates lie between the graphs of
y=(1/2)|x-3|-2 and y=3?
I am just confused on everything here. If you have time, can you help me. Thankyou... Kelly

Answer by stanbon(75887)

How many points with integer coordinates lie between the graphs of
y=(1/2)|x-3|-2 and y=3?
--------------------
y = (1/2)(x-3)-2 or y = (1/2)(3-x)-2
y = (1/2)x - (7/2) or y = (3/2) - (1/2)x -2
y = (1/2)x - (7/2) or y = (-1/2)x - (1/2)
------------------------------------------------

------------
The number of points between the three lines can be couned by you.
======================
Cheers,
Stan H.